[A number has a ones digit that is half the tens digit. The tens digit is a 8. The hundreds digit is twice the ones digit. The thousands digit is 1 lees then the tens digit. What is the number?]
Found 2 solutions by TimothyLamb, MathTherapy:
Answer by TimothyLamb(4379)

x = ones
y = tens
z = hundreds
u = thousands
---
x = y/2
y = 8
z = 2x
u = y - 1
---
put the system of linear equations into standard form
---
2x - y = 0
y = 8
2x - z = 0
y - u = 1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 4
y = 8
z = 8
u = 7
the number = 7884
---
by substitution:
---
y = 8
x = y/2
x = 4
z = 2x
z = 8
u = y - 1
u = 7
---
solution:
x = 4
y = 8
z = 8
u = 7
the number = 7884
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by MathTherapy(10551)


[A number has a ones digit that is half the tens digit. The tens digit is a 8. The hundreds digit is twice the ones digit. The thousands digit is 1 lees then the tens digit. What is the number?]

Tens digit: 8
Units digit: ½ of tens digit____½(8), or 4
Hundreds digit: 2(4), or 8
Thousands digit: 8 – 1, or 7
Number: 7,884